3. You are at a bank working to secure a loan. The advisor indicates the loan has an APR (simple interest fee) of 10-percent per year, but you read the fine print to see the loan actually has an APY (compound interest fee) of 11.5 percent per year. You use your understanding of simple and compound interest to determine the cash differential owed on a $30,000 loan over five years. The differential solution will be closest to?
a. $3,315 c. $6,700
b. $4,800 d. $8,150
Answer #1

The loan amount is 30000 dollars and the time period for repayment is 5 years. If there is a simple interest of 10% per year, then total accumulated interest amount is 10%*30000*5 = 15000 dollars.

Total interest if we use compound interest is arrived at

30000*(1+11.5%)^5 - 30000 = 21700.60

Difference in the interest amount is 21700 - 15000 = 6700

Select option C.